Determine the volume of cement slurry required to set in the annulus of 20inch conductor pipe, and 107/8- inch casing. Total length of 20inch conductor pipe is 80feet for 107/8 inch section

1
Expert's answer
2021-11-04T00:41:10-0400


Annular Volume bbls = (Dh^2 – OD^2) / 1029.4 x Length in feet, where Dh is the hole diameter.


(20^2 - 13.37^2)/1029.4 = 0.2149


Slurry volume = o.2149*80ft = 527 ft = 17.19


Answer = 17.19bbl